OpenStudy (anonymous):

Would Appreciate Help with The Following Question: Your piggy bank has a total of 46 coins in it; some are dimes and some are quarters. If you have a total of $7.00, how many quarters and how many dimes do you have? (2 points)

OpenStudy (anonymous):

It's alright if you can't help! Just asking :D

OpenStudy (anonymous):

30 dimes and 16 quarters

OpenStudy (anonymous):

How are you so sure :O

OpenStudy (anonymous):

if you add it back up then itll make the right amount ahah.

OpenStudy (anonymous):

check yo work ;) substitute it back in and itll make 7 dollars!

OpenStudy (anonymous):

Thanks Smilekbug, hehe I'm like so lost >.<

OpenStudy (anonymous):

x = dimes, y=quarters. .1x+.25y=7.00 and x+y=46, substitution...y=46-x, sub that back into .1x+.25y=7 and you get x=30. 30+y=46, y=16

OpenStudy (anonymous):

i just got showed up

OpenStudy (anonymous):

That's helpful since I like seeing work, so if I face the type of problem again I know how to work it out meow. <3

OpenStudy (anonymous):

and nawh smileyy, you helped tooo :*

OpenStudy (anonymous):

c:

OpenStudy (anonymous):

omg I'm so hungry and why do I keep typing on here ugh so bored.